Circular No. 7901

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002cv IN NGC 3190 V. Larionov and A. Arkharov, Pulkovo; A. Caratti o Garatti and A. Di

Paola, Rome; and M. Dolci, Teramo, report the discovery of an apparent supernova (J = 15.7, H = 15.3, K = 14.4) on J, H, and K images taken with the 1.1-m AZT-24 telescope at Campo Imperatore on May 13.7–13.9 UT, while monitoring SN 2002bo (cf. IAUC 7847). The new object was confirmed at I = 17.7 by the same group (plus E. Di Carlo, A. Pietrinferni, S. Gagliardi, E. Brocato, A. Tornambé, and F. Bernardi) on two I-band images obtained at the 0.72-m Teramo TNT telescope on May 14. SN 2002cv is located at α ≃ 10^{h}18^{m}03^{s}.68, δ ≃ +21^{o}50′06′′.2 (equinox 2000.0), which is 18′′ west and 10′′ north of the nucleus of NGC 3190; the new object appears to be embedded in a region of high absorption, as 30- and 40-min V -band images show no star there (limiting mag 20). Additional photometry: Mar. 28, I [17: (TNT image); Apr. 29, May 5, and 6, J [18: (AZT-24 images); May 9, J > 18 (very faint star visible).

T. Matheson, K. Stanek, S. Jha, P. Challis, and R. Kirshner, Harvard-Smithsonian Center for Astrophysics, report that images of SN 2002cv, obtained by Stanek on May 15.2 UT with the MMT 6.5-m telescope (+ Minicam), confirm the presence of the new object (at R ∼ 19.0, I ∼ 16.6; calibration provided by W. Li) at position end figures 03^{s}.7, 06′′.1.

M. Turatto, Osservatorio Astronomico di Padova, writes that S. Desidera, S. Benetti, G. Altavilla, A. Pastorello, M. Riello, E. Cappellaro, and he obtained a spectrogram (range 550–1000 nm, resolution 5 nm) of SN 2002cv with the 1.8-m Mt. Ekar telescope (+ AFOSC) on May 15.9 UT, showing a very red spectrum with no signal blueward of 600 nm. In addition to telluric absorptions, no feature is clearly visible; this is consitent with a very reddened spectrum of a core-collapse supernova soon after burst.

COMET C/2002 C1 (IKEYA-ZHANG)

Ephemeris extension to IAUC 7876 (elements MPC 45335):

2002TT α_{2000} δ_{2000} Δ r ϵ β m_{1}
May 6 17^{h}59.^{m}68 +53^{o}10.^{′}2 0.416 1.153 99.^{o}5 59.^{o}7 4.6
16 16 37.19 +38 31.6 0.479 1.319 120.4 41.4 5.4
26 15 57.47 +25 30.4 0.591 1.482 133.3 29.8 6.2
June 5 15 37.05 +15 45.2 0.741 1.640 137.7 24.6 7.1
15 15 26.71 + 8 38.3 0.922 1.794 135.7 23.3 7.9
25 15 22.47 + 3 20.5 1.125 1.945 130.4 23.5 8.6
2002 May 16  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green